A study for the treatment of patients with​ HIV-1 was a​ randomized, controlled,​ double-blind study that compared the effectiveness of​ ritonavir-boosted darunavir​ (rbd), the drug currently used to treat​ HIV-1, with​ dorovirine, a newly developed drug. Of the 375 subjects taking​ ritonavir-boosted darunavir, 307 achieved a positive result. Of the 375 subjects taking​ dorovirine, 321 achieved a positive result. Complete parts​ (a) and​ (b).

Find the sample percentage of subjects who achieved a positive outcome in each group.

The sample percentage of subjects who achieve a positive outcome with​ ritonavir-boosted darunavir is ____​%.

The sample percentage of subjects who achieve a positive outcome with dorovirine is ____%.
Answer #1

P = X / n = 307/375 = 0.8187

The sample percentage of subjects who achieve a positive outcome with​ ritonavir-boosted darunavir is 0.8187 * 100 = 81.87 ​%.

P = X / n = 321/375 = 0.856

The sample percentage of subjects who achieve a positive outcome with dorovirine is 0.8560 * 100 = 85.60 %.